How do I solve -5/3v+4/3=4/7-5/7 ? And can you explain it? thank you
Answer by stanbon(75887)

solve (-5/3)v + (4/3) = (4/7) - (5/7)
-----------
The least common multiple of the denominators is 21
Multiply thru by 21 to get:
-35v + 28 = 12 - 15
-35v = -3-38
-35v = -41
v = 41/35
==============
Cheers,
Stan H.